OpenStudy (anonymous):

State any excluded values. 2(a-1)/3(a+1)

OpenStudy (johnweldon1993):

When a denominator is 0...we have an undefined function because we cannot divide by 0 So all you need to do is find out when \[\large 3(a + 1) = 0\] what value of 'a' makes that equation = 0?

OpenStudy (anonymous):

idk negative something?

OpenStudy (johnweldon1993):

Right... 3(a + 1) = 0 Well 3 is being multiplied by those parenthesis...any we know we want 0...well anything times 0 = 0 ...so lets make those parenthesis = 0 (a + 1) = 0 Must mean a = -1 right?

OpenStudy (anonymous):

3(a+1)=0 ;so the ans:a=-1

OpenStudy (anonymous):

So on my paper do i just write: 3(a+1)=0 ;so the ans:a=-1